# A Bayesian Generative Modeling Approach for Arbitrary Conditional Inference

### v1

@@ 5 Conclusion
We also established theoretical guarantees for the BGM framework, including consistency and finite-sample risk control under mild regularity conditions, showing that generative modeling can be competitive with methods that are tailored directly to regression. Empirically, across a range of simulated and real-data experiments, BGM consistently delivered accurate point predictions and well-calibrated uncertainty, outperforming strong discriminative and conformal prediction baselines in various settings. The data imputation experiments on MNIST dataset further illustrated that a single trained BGM can serve as a versatile engine for data imputation under arbitrary missing patterns, benefiting downstream tasks, such as classification.
@@
The Linear Regression was implemented using the LinearRegression class from the scikit-learn library (Pedregosa et al., 2011) with default parameters. We employed random forest regression as a flexible ensemble-based nonparametric baseline (Breiman, 2001), implemented using the RandomForestRegressor class from the scikit-learn library with default hyperparameters. Gradient boosted decision trees were implemented using the XGBoost Python package (Chen, 2016). We trained an XGBRegressor with squared error loss, using 500 boosting iterations, a learning rate of 0.05, and a maximum tree depth of 4. Subsampling was applied to both observations and features, with subsample and column subsample ratios set to 0.8. These settings follow common practice to balance predictive accuracy and regularization. XGBoost represents a strong ensemble-based machine learning baseline that sequentially improves predictions by fitting trees to residual errors.
@@
All conformal prediction methods were implemented using the official github repository (https://github.com/LeyingGuan/LCPexperiments) for localized conformal prediction (LCP) (Guan, 2023). The vanilla CP constructs prediction intervals based on absolute residuals from a fitted point predictor using a held-out calibration set. The nonconformity score is defined as the absolute prediction error. LW-CP normalizes residuals by an estimated local noise level to account for heteroscedasticity (Lei et al., 2018). Local variance estimates are obtained via nearest-neighbor smoothing in the predictor space. QR-CP constructs prediction intervals by conformalizing estimated conditional quantiles (Romano et al., 2019). Lower and upper quantile regressions were trained using the neural network predictor, and conformal scores were formed based on quantile violations. LWQR-CP further adjusts quantile-based conformal scores using local variability estimates, improving adaptivity in heterogeneous settings. Localized CP (Guan, 2023) extends standard CP by weighting calibration samples according to their similarity to the test point. Similarity was measured using Euclidean distance in the learned feature representation of the neural network predictor. Applying localization to the above score constructions yields three additional baselines: LW-LCP, QR-LCP, and LWQR-LCP. These methods combine the respective nonconformity scores with localized weighting schemes to improve empirical adaptivity.
@@
Missing values from MNIST testing set were imputed using the empirical mean of each variable computed from the observed entries. This mean imputation baseline provides a simple and fast reference method and was implemented using SimpleImputer class from the scikit-learn library. MICE (Van Buuren and Groothuis-Oudshoorn, 2011) method teratively imputes missing values using conditional regression models for each variable. We implemented MICE using IterativeImputer class from the scikit-learn library with default settings.

### v2

@@ 5 Conclusion
We also established theoretical guarantees for the BGM framework, including consistency and asymptotic risk control under mild regularity conditions, showing that generative modeling can be competitive with methods that are tailored directly to regression. Empirically, across a range of simulated and real-data experiments, BGM consistently delivered accurate point predictions and uncertainty quantification, outperforming strong discriminative and conformal prediction baselines in various settings. The data imputation experiments on MNIST dataset further illustrated that a single trained BGM can serve as a versatile engine for data imputation under arbitrary missing patterns, benefiting downstream tasks, such as classification.
@@
6 Acknowledgement
The authors are grateful to Faming Liang, Xiaotong Shen, and Leying Guan for their constructive feedback.
7 Funding
Q.L. was partially supported by NIH grant R00 HG013661. W.H.W. was partially supported by NSF-DMS 2310788.
8 Disclosure Statement
The authors report there are no competing interests to declare.
@@ 8 Disclosure Statement
The Linear Regression was implemented using the LinearRegression class from the scikit-learn library (Pedregosa et al. 2011) with default parameters. We employed random forest regression as a flexible ensemble-based nonparametric baseline (Breiman 2001), implemented using the RandomForestRegressor class from the scikit-learn library with default hyperparameters. Gradient boosted decision trees were implemented using the XGBoost Python package (Chen 2016). We trained an XGBRegressor with squared error loss, using 500 boosting iterations, a learning rate of 0.05, and a maximum tree depth of 4. Subsampling was applied to both observations and features, with subsample and column subsample ratios set to 0.8. These settings follow common practice to balance predictive accuracy and regularization. XGBoost represents a strong ensemble-based machine learning baseline that sequentially improves predictions by fitting trees to residual errors.
@@
All conformal prediction methods were implemented using the official github repository (https://github.com/LeyingGuan/LCPexperiments) for localized conformal prediction (LCP) (Guan 2023). The vanilla CP constructs prediction intervals based on absolute residuals from a fitted point predictor using a held-out calibration set. The nonconformity score is defined as the absolute prediction error. LW-CP normalizes residuals by an estimated local noise level to account for heteroscedasticity (Lei et al. 2018). Local variance estimates are obtained via nearest-neighbor smoothing in the predictor space. QR-CP constructs prediction intervals by conformalizing estimated conditional quantiles (Romano et al. 2019). Lower and upper quantile regressions were trained using the neural network predictor, and conformal scores were formed based on quantile violations. LWQR-CP further adjusts quantile-based conformal scores using local variability estimates, improving adaptivity in heterogeneous settings. Localized CP (Guan 2023) extends standard CP by weighting calibration samples according to their similarity to the test point. Similarity was measured using Euclidean distance in the learned feature representation of the neural network predictor. Applying localization to the above score constructions yields three additional baselines: LW-LCP, QR-LCP, and LWQR-LCP. These methods combine the respective nonconformity scores with localized weighting schemes to improve empirical adaptivity.
@@
Missing values from MNIST testing set were imputed using the empirical mean of each variable computed from the observed entries. This mean imputation baseline provides a simple and fast reference method and was implemented using SimpleImputer class from the scikit-learn library. MICE (Van Buuren and Groothuis-Oudshoorn 2011) method teratively imputes missing values using conditional regression models for each variable. We implemented MICE using IterativeImputer class from the scikit-learn library with default settings.
Supplementary Tables
Table S1: Comparison of interval estimation performance across BGM and VAEAC. Metrics include Pearson correlation (PCC), Spearman correlation (SCC), empirical marginal coverage, and average prediction interval length (ave.PI) at different dimensions pp. The experimental setting is the same as Table 2. VAEAC failed to capture meaningful uncertainty based on the multiple imputations.
$p=50$ $p=100$ $p=300$
Metric VAEAC BGM VAEAC BGM VAEAC BGM
[…]
Coverage 0.050 0.944 0.056 0.950 0.024 0.966
ave.PI 0.043 1.450 0.060 1.576 0.038 1.694
  • Note: Coverage values closest to the nominal level 0.95 are highlighted in bold. Average interval lengths closest to the ground truth are highlighted in bold.